# A Certified Unlearning Approach without Access to Source Data

### v1

@@ 1 Introduction
This problem is related to zero-shot unlearning, where no information about the true dataset is available during forgetting, beyond the model. While several relevant methods (Foster et al., 2024; Chundawat et al., 2023; Cha et al., 2023) have been proposed, no theoretical guarantees exist.
@@ 2 Related Works
Source-Free Unlearning. Zero-shot machine unlearning (Chundawat et al., 2023) removes data influence using only model weights and the forget set, using noise-based error maximization and gated knowledge transfer. Another method (Cha et al., 2023) employs adversarial sample generation to preserve decision boundaries while applying gradient ascent on forget data. JiT unlearning (Foster et al., 2024) fine-tunes models with perturbed samples to reduce reliance on forget instances. Bonato et al. (Bonato et al., 2025) modify feature vectors to align forgotten data with the nearest incorrect class, using a surrogate dataset in source-free settings. Despite some recent work in zero-shot/source-free unlearning, formal certified guarantees remain an open problem. We address this by ensuring such guarantees using a surrogate dataset in the absence of source data.
@@ 4.2 From Theory to Practice
[…]
@@ B.1 Certified Unlearning Using Newton Updates with Exact Data Samples
Suppose the loss function $\mathcal{L}$ satisfies 4.1. Let $\boldsymbol{w}_{r}$ denote the model approximated using the exact dataset $\mathcal{D}$ of size $n_{1}$ and the forget set $\mathcal{D}_{u}$ of size $m$. Similarly, let $\widehat{\boldsymbol{w}}_{r}$ denote the model approximated using the surrogate dataset $\mathcal{D}_{s}$ of size $n_{2}$. Also, assume that the $n_{1}$ and $n_{2}$ are sufficiently large and $n_{i}\geq\frac{m\beta}{\alpha}$ where $i\in\{1,2\}$ Then, the norm of the difference between the approximated models is upper bounded as follows:
@@
Consider a loss function $\mathcal{L}$ satisfying 4.1, and a surrogate dataset $\mathcal{D}_{s}$ with $n_{2}$ samples drawn from a distribution $\nu$, to mimic the true dataset $\mathcal{D}$ with $n_{1}$ drawn from a distribution $\rho$, over the support set $\mathcal{X}\times\mathcal{Y}$. Define the retrained model over the retained samples as $\boldsymbol{w}^{\ast}_{r}$ and the model achieved after unlearning as $\widehat{\boldsymbol{w}}_{r}$. Also, assume that the $n_{1}$ and $n_{2}$ are sufficiently large and $n_{i}\geq\frac{m\beta}{\alpha}$ where $i\in\{1,2\}$ Then, the following upper bound holds,
@@ Appendix D Approximating Kullbeck-Leiber Distance
[…]
@@
where $\varepsilon\sim\mathcal{N}(0,\mu)$ is Gaussian noise, and $\mu$ represents the step size or learning rate. The negative gradient term $-\frac{\partial E(\boldsymbol{x})}{\partial\boldsymbol{x}}$ directs the samples toward regions of lower energy (higher likelihood), while the Gaussian noise ensures exploration of the input space to avoid convergence to local minima. The process begins with initializing $\boldsymbol{x}_{0}$ from a prior distribution over the input space, which is often chosen to be uniform for simplicity and generality.
@@ Appendix F Additional Synthetic and Real-World Dataset Experiments
Original 78.2$\pm$0.274.0$\pm$0.378.2$\pm$0.278.6$\pm$0.447.6$\pm$0.2%
Retrain 77.2$\pm$0.371.8$\pm$0.277.7$\pm$0.573.2$\pm$0.447.4$\pm$0.210$\pm$1
Unlearn (+) 77.4$\pm$0.672.1$\pm$0.776.9$\pm$0.874.1$\pm$0.347.5$\pm$0.410$\pm$2 0.02
0.02 Unlearn (-) 77.5$\pm$0.272.3$\pm$0.577.8$\pm$0.174.5$\pm$0.249.1$\pm$0.49$\pm$2 0.23 0.21$\pm$0.12
0.04 Unlearn (-) 77.4$\pm$0.672.4$\pm$0.377.2$\pm$0.474.3$\pm$0.349.2$\pm$0.210$\pm$3 0.31 0.35$\pm$0.23
0.06 Unlearn (-) 77.4$\pm$0.572.4$\pm$0.677.5$\pm$0.274.2$\pm$0.548.7$\pm$0.711$\pm$2 0.37 0.38$\pm$0.13
0.08 Unlearn (-) 77.5$\pm$0.772.3$\pm$0.277.4$\pm$0.774.2$\pm$0.348.1$\pm$0.59$\pm$1 0.4 0.39$\pm$0.15
0.1 Unlearn (-) 77.4$\pm$0.472.3$\pm$0.177.7$\pm$0.474.1$\pm$0.948.2$\pm$0.212$\pm$3 0.41 0.41$\pm$0.08
@@
13 Original 85.9$\pm$0.873.3$\pm$0.185.4$\pm$0.584.6$\pm$0.2%
Retrain 86.7$\pm$0.273.7$\pm$0.887.7$\pm$0.575.6$\pm$0.251.5$\pm$0.770$\pm$1
Unlearn (+) 84.1$\pm$0.971.7$\pm$0.485.2$\pm$0.673.2$\pm$0.751.4$\pm$0.415$\pm$3 0.02
Unlearn (-) 84.7$\pm$0.472.2$\pm$0.985.3$\pm$0.773.6$\pm$0.551.8$\pm$0.315$\pm$2 0.49$\pm$0.03 0.51$\pm$0.04
36 Original 85.5$\pm$0.176.2$\pm$0.785.6$\pm$0.883.1$\pm$0.3%
Retrain 84.1$\pm$0.475.1$\pm$0.886.1$\pm$0.971.5$\pm$0.550.3$\pm$0.120$\pm$1
Unlearn (+) 84.9$\pm$0.175.6$\pm$0.185.1$\pm$0.171.5$\pm$0.350.8$\pm$0.518$\pm$1 0.02
Unlearn (-) 84.9$\pm$0.875.3$\pm$0.985.1$\pm$0.472.4$\pm$0.850.2$\pm$0.819$\pm$3 0.43$\pm$0.02 0.41$\pm$0.02
100 Original 84.7$\pm$0.571.3$\pm$0.384.9$\pm$0.483.6$\pm$0.5%
Retrain 82.9$\pm$0.376.0$\pm$0.284.2$\pm$0.871.1$\pm$0.552.5$\pm$0.419$\pm$2
Unlearn (+) 83.8$\pm$0.575.7$\pm$0.785.1$\pm$0.172.2$\pm$0.852.0$\pm$0.421$\pm$1 0.02
Unlearn (-) 83.7$\pm$0.475.6$\pm$0.385.0$\pm$0.772.0$\pm$0.651.9$\pm$0.216$\pm$3 0.25$\pm$0.02 0.31$\pm$0.07

### v2

@@ 1 Introduction
This problem is related to zero-shot unlearning, where no information about the true dataset is available during forgetting, beyond the model. While several relevant methods (Foster et al., 2024; Chundawat et al., 2023; Cha et al., 2023; Ahmed et al., 2025) have been proposed, no theoretical guarantees on certified unlearning exist.
@@ 2 Related Works
Source-Free Unlearning. Zero-shot machine unlearning (Chundawat et al., 2023) removes data influence using only model weights and the forget set, using noise-based error maximization and gated knowledge transfer. Another method (Cha et al., 2023) employs adversarial sample generation to preserve decision boundaries while applying gradient ascent on forget data. JiT unlearning (Foster et al., 2024) fine-tunes models with perturbed samples to reduce reliance on forget instances. Bonato et al. (Bonato et al., 2025) modify feature vectors to align forgotten data with the nearest incorrect class, using a surrogate dataset in source-free settings. In a concurrent work, Ahmed et al. (Ahmed et al., 2025) approximate the unlearning update by collecting perturbed points around local minima and utilizing only the forget set. This work provides theoretical bounds justifying the methodology, but do not consider formal certified guarantees. Despite some recent work in zero-shot/source-free unlearning, formal certified guarantees remain an open problem. We address this by ensuring such guarantees using a surrogate dataset in the absence of source data.
@@ 4.2 From Theory to Practice
$\begin{aligned} &\mathrm{KL}(\nu\ \|\ \rho)\approx\frac{1}{n}\sum_{(\boldsymbol{x},y)\in\mathcal{D}_{s}}\log{\frac{f(\tilde{\boldsymbol{w}}^{\ast},\boldsymbol{x})_{y}}{f(\boldsymbol{w}^{\ast},\boldsymbol{x})_{y}}}\\ &+\sup_{T}\Bigg{(}\frac{1}{n}\sum_{i=1}^{n}T(\boldsymbol{x}_{i})-\log\left(\frac{1}{k}\sum_{j=1}^{k}\exp(T(\widehat{\boldsymbol{x}}_{j}))\right)\Bigg{)}\end{aligned}$
@@ B.1 Certified Unlearning Using Newton Updates with Exact Data Samples
Suppose the loss function $\mathcal{L}$ satisfies 4.1. Let $\boldsymbol{w}_{r}$ denote the model approximated using the exact dataset $\mathcal{D}$ of size $n_{1}$ and the forget set $\mathcal{D}_{u}$ of size $m$. Similarly, let $\widehat{\boldsymbol{w}}_{r}$ denote the model approximated using the surrogate dataset $\mathcal{D}_{s}$ of size $n_{2}$. Also, assume that the $n_{1}$ and $n_{2}$ are sufficiently large and $n_{i}\geq\frac{m\beta}{\alpha}$ where $i\in\{1,2\}$ . Then, the norm of the difference between the approximated models is upper bounded as follows:
@@
Consider a loss function $\mathcal{L}$ satisfying 4.1, and a surrogate dataset $\mathcal{D}_{s}$ with $n_{2}$ samples drawn from a distribution $\nu$, to mimic the true dataset $\mathcal{D}$ with $n_{1}$ drawn from a distribution $\rho$, over the support set $\mathcal{X}\times\mathcal{Y}$. Define the retrained model over the retained samples as $\boldsymbol{w}^{\ast}_{r}$ and the model achieved after unlearning as $\widehat{\boldsymbol{w}}_{r}$. Also, assume that the $n_{1}$ and $n_{2}$ are sufficiently large and $n_{i}\geq\frac{m\beta}{\alpha}$ where $i\in\{1,2\}$ . Then, the following upper bound holds,
@@ Appendix D Approximating Kullbeck-Leiber Distance
where $f(\boldsymbol{w}^{\ast},\boldsymbol{x})_{y}$ corresponds to the logit score (or unnormalized log-probability) for the label $y$ given the input $\boldsymbol{x}$ under the model $\boldsymbol{w}^{\ast}$. The summation is taken over the support of the output label space $\mathcal{Y}$. The normalization constant $Z=\int\exp(-E(\boldsymbol{x}))\,d\boldsymbol{x}$ ensures that $\rho(\boldsymbol{x})$ is a valid probability distribution. However, evaluating $Z$ is computationally intractable due to the high-dimensional integral over the input space.
@@
where $\varepsilon\sim\mathcal{N}(0,\mu)$ is Gaussian noise, and $\mu$ represents the step size or learning rate. The negative gradient term $\frac{\partial E(\boldsymbol{x})}{\partial\boldsymbol{x}}$ directs the samples toward regions of lower energy (higher likelihood), while the Gaussian noise ensures exploration of the input space to avoid convergence to local minima. The process begins with initializing $\boldsymbol{x}_{0}$ from a prior distribution over the input space, which is often chosen to be uniform for simplicity and generality.
@@ Appendix F Additional Synthetic and Real-World Dataset Experiments
[…]
@@
[…]